### v1

# Design of AI-Powered Tool for Self-Regulation Support in Programming Education

Abstract.
Large Language Model (LLM) tools have demonstrated their potential to deliver high-quality assistance by providing instant, personalized feedback that is crucial for effective programming education. However, many of these tools operate independently from institutional Learning Management Systems, which creates a significant disconnect. This isolation limits the ability to leverage learning materials and exercise context for generating tailored, context-aware feedback. Furthermore, previous research on self-regulated learning and LLM support mainly focused on knowledge acquisition, not the development of important self-regulation skills. To address these challenges, we developed CodeRunner Agent, an LLM-based programming assistant that integrates the CodeRunner, a student-submitted code executing and automated grading plugin in Moodle. CodeRunner Agent empowers educators to customize AI-generated feedback by incorporating detailed context from lecture materials, programming questions, student answers, and execution results. Additionally, it enhances students’ self-regulated learning by providing strategy-based AI responses. This integrated, context-aware, and skill-focused approach offers promising avenues for data-driven improvements in programming education.
LLM-powered tools, Programming education, Self-regulated learning, Personalized feedback, CodeRunner
copyright: noneconference: Augmented Educators and AI: Shaping the Future of Human-AI Collaboration in Learning on CHI 2025 Workshop; April 26,2025; Yokohama, JAPANccs: Applied computing Computer-assisted instructionccs: Applied computing → E-learning
© This paper was adapted for the CHI 2025 Workshop on Augmented Educators and AI: Shaping the Future of Human and AI Cooperation in Learning, held in Yokohama, Japan on April 26, 2025. This work is licensed under the Creative Commons Attribution 4.0 International License (CC BY 4.0).
1. Introduction
As programming education attracts an ever-growing number of students, providing immediate, individualized assistance during challenging moments has become increasingly difficult (Chiu, 2024; Yilmaz and Yilmaz, 2023b; Ma et al., 2024a). However, the reality in many educational institutions is far from ideal. Traditional support methods, such as scheduled office hours, are often limited, and in-person help can be both time-consuming and labor-intensive. This situation frequently impedes the delivery of the tailored feedback students need to fully comprehend complex programming concepts. Consequently, there is an urgent need for scalable, equitable, and user-friendly solutions that can offer instant support in programming education (Tian et al., [n. d.]; Yilmaz and Yilmaz, 2023a; Pankiewicz and Baker, 2023). In this context, the advent of LLMs presents a promising alternative. A variety of LLM-powered programming assistants have been developed to provide timely coding guidance and suggestions, potentially transforming the landscape of programming education.
While these advancements present exciting opportunities for personalized learning and efficient problem-solving, they also raise critical questions about the role and effectiveness of LLM-powered assistants in truly enhancing student learning. One major concern is that students may become overly reliant on these tools, potentially hindering the development of self-regulated learning (SRL) skills and problem-solving skills (Anagnostopoulos, 2023; Rajala et al., 2023; Humble et al., 2023). Research indicates that the ease of obtaining direct answers can foster a superficial understanding, as students may bypass the rigorous process of working through challenges independently (Skjuve et al., 2023; Tlili et al., 2023). Unfortunately, most existing LLM-powered programming tools fail to address this issue and ignore the implementation of pedagogically sound guardrails to restrict the LLM’s capacity to guide students to engage in self-regulated learning and use AI as a supportive aid (Sun et al., 2024).
Another significant issue is that many LLM-based tools operate independently of institutional Learning Management Systems (LMS) like Moodle. This separation creates a disconnect between the tool and the broader educational context, limiting the opportunity to leverage rich assignment details and course materials for more tailored, context-aware feedback. For example, Ma et al. (Ma et al., 2024a) found that when LLMs are not aligned with a student’s specific curriculum, they can generate advanced or off-topic responses that stray from the intended teaching scope, ultimately detracting from the learning process. Without seamless integration, educators struggle to track how students interact with LLM-generated feedback over time, making it difficult to assess the true impact on learning outcomes (Ma et al., 2024b). Therefore, incorporating LLM-based tools within the LMS environment is essential to ensure that the feedback provided is both relevant and aligned with course objectives. Additionally, the effectiveness of AI-generated feedback is highly sensitive to the context of the learning tasks. The lecture materials and exercises in programming education are highly important contextual information; however, they are not well connected in current LLM-based scaffolding research (Kazemitabaar et al., 2024). The Learning Analytics (LA) technique has to enhance the quality of LLM-based feedback by utilizing the behavioral data in context between learners, lecture materials, and exercise solving (Kaur and Chahal, 2024).
To address these limitations, we developed CodeRunner Agent, a LLM-based programming assistant that seamlessly extends CodeRunner, a free and open-source plug-in for Moodle designed to execute and assess student-submitted code. CodeRunner Agent is designed to meet the needs of both students and educators. It allowed to leverage the comprehensive context available within an LMS environment from learning logs and enhance students’ self-regulated learning in introductory programming education. This enables the delivery of customized feedback that not just give the correct answer but enhances students’ self-regulation and problem solving skills. Beyond timely and strategy-based feedback, the CodeRunner Agent continuously tracks AI responses and subsequent modifications in students’ code. This capability provides educators and researchers with valuable insights into learning trajectories and the overall effectiveness of AI-assisted instruction. By collecting and analyzing data on how students interact with and adapt to AI-generated feedback, our approach paves the way for data-driven improvements in programming education. In summary, CodeRunner Agent represents a significant step forward in integrating AI tools within institutional LMS environments to enhance programming education. By addressing the challenges of scalability, contextual feedback, and skill enhancement, our work offers promising avenues for improving student achievements and deepening our understanding of AI’s role in modern education.
2. Related Work
As LLMs become increasingly pervasive, educational researchers are examining their potential to generate educational content, boost student engagement, and personalize learning experiences. This is particularly relevant in programming education, where the adoption of such tools is prompting a reevaluation of traditional teaching methods (Shoufan, 2023; Kasneci et al., 2023).
Recent studies have primarily focused on assessing LLMs’ capabilities in programming tasks, ranging from code generation and program repair to code explanation and code summarization (Rajala et al., 2023). For instance, Finnie-Ansley et al. (Finnie-Ansley et al., 2023) demonstrated that OpenAI Codex outperforms most students on code-writing questions in both CS1 and CS2 exams. In a similar vein, Savelka et al. (Savelka et al., 2023) evaluated GPT-3 and GPT-4 on programming exercises across three Python courses, revealing that these models progressed from failing typical assessments to passing courses without human intervention. Furthermore, Sarsa et al. (Sarsa et al., 2022) examined programming exercises generated by OpenAI Codex, assessing their novelty, plausibility, and readiness, and highlighted the potential for these models to create effective coding assignments. Recent work by Phung et al. (Phung et al., 2023) systematically compared GPT models with human tutors, finding that they approach human-level performance in both Python programming tasks and the resolution of real-world buggy programs. Additionally, ChatGPT has proven effective in providing feedback on programming assignments and aiding students in applying theoretical knowledge practically. Prior research has underscored the model’s capacity to generate personalized feedback that students rate positively (Pankiewicz and Baker, 2023).
While LLMs hold significant promise for enhancing programming education, both students and researchers have expressed concerns about their direct use. One major worry is that students might become overly reliant on LLMs, potentially stunting the development of SRL skills. Additionally, many students struggle with formulating effective prompts, often resulting in feedback that fails to meet their learning needs. In response, researchers have increasingly developed specialized LLM-based tools that address these issues. For instance, Kazemitabaar et al. (Kazemitabaar et al., 2023) developed Coding Steps, which leverages LLM-based code generators to support beginners in introductory programming courses. Similarly, Lifton et al. (Liffiton et al., 2023) introduced the CodeHelp tool, designed to assist students while incorporating guardrails that prevent the tool from directly revealing complete solutions. With CodeHelp, students can input a free-form question along with their code and, optionally, an error message, ensuring that the feedback remains contextual and instructive. In another example, CodeAid (Kazemitabaar et al., 2024) offers a range of input templates and interactive response formats tailored to diverse student needs. It employs scaffolding techniques, such as interactive pseudo-code and detailed code annotations, to guide students from grasping fundamental programming concepts to independently writing and debugging their code. These innovations collectively illustrate the emerging trend of developing LLM-based tools that not only harness the power of AI but also promote deeper learning and independence among students.
Despite the impressive achievements of these tools in enhancing programming education, they often operate in isolation, lacking the integration of various functionalities and failing to connect with Integrated Development Environments (IDE) and Learning Management Systems (LMS). This fragmented approach does not adequately meet the needs of educators and students, thereby hindering widespread adoption and scalability. To address these limitations, we developed CodeRunner Agent, a comprehensive solution that seamlessly integrates LLM-powered assistance with both IDEs and LMS platforms. By unifying these functionalities, our system offers a more cohesive and effective learning environment, ensuring that both instructors and learners have access to timely, context-aware support on a large scale.
3. CodeRunner Agent
3.1. Overview of CodeRunner Agent
The framework of CodeRunner Agent is shown in Figure 1. It is developed and integrated to scaffold programming education with an LLM-based assistant in the Moodle LMS. The framework contains a lecture viewer, a CodeRunner plugin and CodeRunner Agent.
Figure 1. Overview of Programming Support Environment
The lecture viewer is provided to deliver the lecture slides by instructors and access the lecture slides by learners inside and outside of class. The operations of the lecture viewer are recorded in the form of Experience API (or xAPI) statements. Then the xAPI statements are stored in the Learning Record Store (LRS). The examples of the operation logs are accessing time and accessing frequency.
CodeRunner (Lobb and Harlow, 2016) is a free, open-source plugin and it can be imported to Moodle LMS. Learners can write programming codes to solve programming problems and receive automated grades by running it in a series of tests. A logging plugin named “Logstore xAPI” (Rotelli et al., 2023) is used to record the CodeRunner results and send them to the LRS. For example, the code for the test, got output, correct status, mark reward will be logged in LRS if learners run the test code once for testing in CodeRunner.
CodeRunner Agent is an AI-powered tool to support learners’ SRL and teachers’ customized designs for LLM-based feedback in programming education. It can be embedded into the Moodle LMS and executed with the CodeRunner plugin. Learners can receive general-purpose and programming-specific regulatory strategy hints from LLMs-based feedback. Instructors can upload learning materials to the context engine of the CodeRunner Agent and customize the contextual parameters of the agent to create tailored instruction. The interactions with the CodeRunner Agent are automatically tracked as xAPI statements and stored in the LRS. For instance, the request type, request time, and exercise ID related to the request will be recorded in LRS if learners send a hint request.
3.2. SRL Support Model in CodeRunner Agent
The SRL scaffoldings in CodeRunner Agent are implemented using a five-phase cycle model, named PPESS (Planning – Program creation – Error correction - Self-monitoring - Self-reflection). The PPESS model is grounded in the well-known Zimmerman’s SRL theory (Zimmerman, 2008) and adopts a conceptual framework for regulating learning in programming (Silva et al., 2024). The model contains five phases: planning, program creation, error correction, self-monitoring, and self-reflection.
The planning phase constitutes the initial stage in code development, encompassing methodologies employed to develop a comprehensive understanding of exercise requirements and identify essential programming elements. The program creation phase represents an implementation stage, wherein learners execute preliminary plans regarding necessary programming components. This phase involves the combination of declarative and procedural knowledge within the programming task, transforming theoretical constructs into functional code. The error correction phase encompasses multiple strategic approaches for addressing coding errors. These include the externalization of meta-cognitive processes to remedy misunderstanding and reflect on limitations, utilization of cognitive strategies for problem identification and solution formulation, and implementation of behavioral regulation techniques. The progress monitoring phase occurs concurrently with the program creation and error correction phases. During this phase, students continuously monitor their progress in completing the task requirements and assess the adequacy of their code regarding programming syntax and standards. Finally, the self-reflection phase, occurring at the end of code development, enables learners to identify areas requiring improvement, maintain motivation throughout their programming, and engage in evaluation regarding performance and behavior for enhancing learning experiences and skill development.
The phase, target SRL strategy, and LLMs support in the CodeRunner Agent are summarized in Table 1. Both general-purpose and programming-specific regulatory strategies are included for SRL scaffoldings in the five-phase PPESS model. To enhance strategy-based programming learning for novice learners, LLM-powered feedback is integrated into the five-phase PPESS model by leveraging LLMs’ capabilities as personalized, adaptive scaffolds.
Table 1. Phase, Target SRL strategy, LLMs support in the CodeRunner Agent.
Phase Strategy LLM-powered feedback
Planning Problem understanding, problem definition, program logic planning LLMs offer the basic knowledge of the exercise’s requirements and suggest planning the program step-by-step using diagrams, pseudocode, or notes.
Program creation Review lecture materials, review previous exercise, code dividing, code commenting LLMs provide the location of required knowledge in lecture materials, supplemental resources related to the exercise, and explanations for the key points.
Error correction Review the exercise statement, utilize test cases, analyze the error message, emotional regulation, help-seeking LLMs provide suggestions for effective error correction and generate hints on fixing syntactic and logical errors without showing the solution directly.
Self-monitoring Check the progress of all exercises, test the program regularly LLMs encourage learners to track their own learning progress regularly.
Self-reflection Achievement self-assessment, effort self-assessment, code review, code refactoring LLMs give evaluations on students’ behavioral process and final performance, motivate students by finding their strength points or the effort they put in, and suggest students to identify their improvement areas.
Figure 2. User Interface of Integrated CodeRunner Agent for Learners
The user interface of the integrated CodeRunner Agent for learners is shown in Figure 2. The dialogue window of the CodeRunner Agent is displayed immediately below the test results of CodeRunner. Learners can select one SRL phase from five phases: planning, program creation, error correction, self-monitoring, and self-reflection. They can choose one request type from two types: general purpose or programming-specific. After sending the hint request, the LLM-based strategy hint will be shown in the response window of the CodeRunner Agent.
3.3. Context Engine in CodeRunner Agent
The CodeRunner Agent’s core intelligence contains two context engines: the Learning Analytics Context Engine (LACE) and the Knowledge Context Engine (KCE). The LACE calculates and aggregates learners’ engagement metrics (i.e., time spent, attempt frequency) and performance metrics (i.e., success rates, error patterns) from LRS data after filtering personal data (i.e., name, gender, and email). The KCE handles both lecture and exercise knowledge bases. The lecture knowledge base manages the key programming concepts and the dependencies between the concepts from lecture materials. On the other hand, the exercise knowledge base categorizes exercises by concept, difficulty, solution, and typical mistakes.
Instructors can upload lecture materials (i.e., concept definition, concept illustration, and annotated examples) and exercises (i.e., problem statement, solution, and test cases) to the context engine. The lecture materials and exercises will be converted to textual knowledge and stored in the knowledge database for future retrieval. More importantly, instructors can update the knowledge bases and customize the parameters of the context engine by using the configuration tools.
4. Discussion and Future work
This study aims to develop an LLM-based programming assistant, CodeRunner Agent, that seamlessly integrates with a lecture viewer and CodeRunner plugin in the Moodle LMS. The design is grounded within Zimmerman’s SRL theory and targets the freshmen students’ programming-specific knowledge acquisition as well as general self-regulation skill development. This study can fill the gap between AI, SRL, and learning analytics (LA) by combining LLM-based scaffolding, self-regulation in programming, and learning log-based contextual feedback.
Future work will be conducted for the interface improvement of the configuration tools and user-friendly workflow in the instructor-agent collaboration. Additionally, the effects of the CodeRunner Agent will be evaluated through short-term pilot studies in an actual class and semester-long experiments in multiple actual classes in university.